OpenStudy (anonymous):

prove 1+ 2^n < 3^n for n>=2

OpenStudy (anonymous):

induction step

OpenStudy (anonymous):

Is it true for \(n=2\) (the base case)? \[1+2^2=5<9=3^2\] Yes. Assume it's true for \(n=k\): \[1+2^k<3^k\] Use this to show it's true for \(n=k+1\). What you want to show: \[(*)~~~~1+2^{k+1}<3^{k+1}\] Rewriting \((*)\), you could have \[1+2(2^k)=\color{red}{1+2^k}+2^k<\color{red}{3^k}+2^k\] The red part is true by the induction hypothesis. We need to establish that \(3^k+2^k\le3^{k+1}\). \[3^{k+1}=3(3^k)=3^k+3^k+3^k\] If you can accept that \(2^k<3^k\) (you might have to prove) for all \(k\ge1\), you're done, since \[\color{red}{3^{k+1}}=3^k+3^k+3^k\color{red}>3^k+2^k+2^k=\color{red}{3^k+2^{k+1}>3^k+2^k>1+2^{k+1}}\]

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

You're welcome